# Tatami Printer: Physical ZKPs for Tatami Puzzles

### v1

@@ Abstract
Tatami puzzles are pencil-and-paper logic puzzles with an objective to partition a rectangular grid into rectangular regions such that no four regions share a corner point, as well as satisfying other constraints. In this paper, we develop a physical card-based protocol called Tatami printer that can help verify solutions of Tatami puzzles. We also use the Tatami printer to construct physical zero-knowledge proof protocols for two such puzzles: Tatamibari and Square Jam. These protocols enable a prover to show a verifier the existence of the puzzles’ solutions without revealing them.
@@ 1 Introduction
Pencil puzzles are logic puzzles that can be solved by writing down the solution on a paper. Examples of them include Sudoku, Numberlink, and Nonogram. Pencil puzzles can be classified into several categories based on their main theme. Puzzles having a theme of partitioning a rectangular grid into multiple regions to satisfy certain rules are called decomposition puzzles.
Many decomposition puzzles, especially the ones originated in Japan, specify that the regions must be rectangular, and also have a common constraint that no four regions can share a corner point. This rule came from the arrangement of Japanese Tatami mats. Hence, these puzzles can be collectively called Tatami puzzles, a subcategory of decomposition puzzles [10].
@@ 1.3 Zero-Knowledge Proof
Suppose Agnes created a difficult pencil puzzle and challenged her friend Brian to solve it. After a while, he could not solve the puzzle and began to doubt whether it actually has a solution. Agnes needs to convince him that her puzzle has a solution without revealing it (which would render the challenge pointless). A zero-knowledge proof (ZKP) makes this seemingly difficult task possible.
@@
In 1991, Goldreich et al. [2] proved that a ZKP exists for every NP problem, implying one can construct a computational ZKP for any NP pencil puzzle via a reduction. Such construction, however, is neither practical nor intuitive as it requires cryptographic primitives. Instead, many researchers focused on developing physical ZKPs using objects found in everyday life such as a deck of playing cards. These card-based protocols have benefits that they do not require computers and also allow external observers to verify that the prover truthfully executes them (which is often a challenging task for digital protocols). They also have didactic values and can be used to teach the concept of ZKP to non-experts.
@@ 1.4 Card-based Zero-Knowledge Proof Protocols
There is a line of work dedicated to construct physical card-based ZKP protocols for specific pencil puzzles, such as ABC End View [13], Goishi Hiroi [13], Heyawake [11], Hitori [11], Juosan [8], Kakuro [9], Masyu [7], Nonogram [12], Numberlink [15], Nurikabe [11], Ripple Effect [16], Shikaku [14], Slitherlink [7], Sudoku [18], Takuzu [8], and Toichika [13].
In 2024, Ruangwises and Iwamoto [17] proposed a more generic protocol called printing protocol, which can help verify solutions of many decomposition puzzles such as Five Cells and Meadows. Their protocol, however, has a major drawback that it cannot check constraints involving a relation between multiple regions, so it cannot be used to verify solutions of Tatami puzzles.
@@ 1.5 Our Contribution
In this paper, we develop a card-based protocol called Tatami printer, which is modified from the printing protocol of Ruangwises and Iwamoto [17].
@@
We use the Tatami printer to construct ZKP protocols for two Tatami puzzles: Tatamibari and Square Jam.
@@ 2.2 Pile-Shifting Shuffle
Given a matrix $M$ of cards, a pile-shifting shuffle [19] rearranges the columns of $M$ by a uniformly random cyclic shift unknown to $V$. It can be implemented by putting all cards in each column into an envelope, and repeatedly picking some envelopes from the bottom and putting them on the top of the pile.
Note that each card in $M$ can be replaced by a stack of cards, and the protocol still works in the same way as long as every stack in the same row has the same number of cards.
@@ 2.3 Chosen Cut Protocol
Note that each card in $C$ can be replaced by a stack of cards, and the protocol still works in the same way as long as every stack has the same number of cards.
@@ 2.4 Printing Protocol of Ruangwises and Iwamoto
Given a $p\times q$ matrix of face-down cards (called a template) and another $p\times q$ matrix of face-down cards of an area from the puzzle grid, a printing protocol [17] verifies that positions in the area corresponding to non-blank cards in the template are initially empty (consisting of all blank cards). Then, the protocol places all non-blank cards from the template at the corresponding positions in the area, replacing the original blank cards (see Fig. 4) without revealing any card to $V$.
@@
Figure 4: Examples of a successful print from a $3\times 5$ template onto a $3\times 5$ area (top) and an unsuccessful print because of overlapping numbers (bottom)
@@
      Reveal the selected card that it is a blank card (otherwise $V$ rejects) and remove it from the stack.
After these steps, all non-blank cards from the template are placed at the corresponding positions in the area. $V$ is also convinced that these positions in the area were initially empty (consisting of all blank cards).
$P$ can repeatedly apply the printing protocol to print each region of the partition according to $P$’s solution onto the puzzle grid. $V$ will be convinced that these regions do not overlap with one another.
However, a major drawback of this protocol is that each region is separately printed, so it cannot check constraints involving a relation between multiple regions. Therefore, it cannot be used to verify solutions of Tatami puzzles.
@@ 3 Our Proposed Protocol: Tatami Printer
Considering the drawback of the printing protocol, we aim to modify the protocol so that it can also check the corner constraint of Tatami puzzles.
@@
On top of each card in a template, we place a $\clubsuit$  if its top-left corner is a corner point of a rectangle the template represents, and a $\heartsuit$  if its top-left corner is not a corner point. Also, on top of each card in the puzzle grid, we place two cards (initially both $\heartsuit$ s) to keep track of the number of rectangles having its top-left corner as their corner point, which is equal to the number of $\clubsuit$ s among them.
@@
When printing a $p\times q$ template onto a $p\times q$ area from the grid, $P$ performs the following steps.
@@
      Reveal the selected card that it is a blank card (otherwise $V$ rejects) and remove it from the stack.
@@
    For the corner counting part, place each card from the template on top of a corresponding two-card stack from the area, creating $pq$ three-card stacks.
@@
      Reveal the selected card that it is a $\heartsuit$  (otherwise $V$ rejects) and remove it from the stack.
After these steps, all non-blank cards from the main part of the template are placed at the corresponding positions in the area. All $\clubsuit$ s from the corner counting part of the template are also added to the corresponding positions in the area. $V$ is convinced that these positions in the area were initially empty (consisting of all blank cards), and that each point in the area is a corner point of at most two rectangles.
@@ 4.1 Main Protocol
Suppose that in $P$’s extended solution, the grid is partitioned into $k$ rectangles $B_{1},B_{2},...,$ $B_{k}$. For each $i=1,2,...,k$, $P$ performs the following steps.
@@
Finally, $P$ reveals all cards from the main part on the cells that contain a symbol in the original Tatamibari puzzle. $V$ verifies that the symbols on the cards match the ones in the cells (otherwise $V$ rejects). $P$ also reveals the main part of all dummy stacks that they are still blank. If all verification steps pass, then $V$ accepts.
@@ 4.2 Proof of Correctness and Security
Hence, we can conclude that $V$ always accept. ∎
@@ 5 ZKP Protocol for Square Jam
In our protocol, $P$ constructs $n$ templates, one for each size of square. Each template has size $(n+1)\times(n+1)$, and the square is placed at the top-left corner of the template. A cell inside the square is represented by a card with a number equal to the side length of that square, while a cell outside the square is represented by a blank card (see Fig. 7).
@@ 5.1 Main Protocol
$P$ also constructs $n$ templates of all $n$ sizes of square and lets $V$ verify that all templates are correct (otherwise $V$ rejects).
Suppose that in $P$’s extended solution, the grid is partitioned into $k$ squares $B_{1},B_{2},...,$ $B_{k}$. For each $i=1,2,...,k$, $P$ performs the following steps.
@@
    Apply the chosen cut protocol to select an $(n+1)\times(n+1)$ area whose top-left corner is the top-left corner of $B_{i}$.
@@
    Apply the chosen cut protocol to select a template of a square with the same size as $B_{i}$.
@@
    Reconstruct a template that has just been used and replenish the pile of templates with it. Let $V$ verify again that all $n$ templates are correct (otherwise $V$ rejects).
Finally, $P$ reveals all cards from the main part on the cells that contain an integer in the original Square Jam puzzle. $V$ verifies that the integers on the cards match the ones in the cells (otherwise $V$ rejects). $P$ also reveals the main part of all dummy stacks that they are still blank. If all verification steps pass, then $V$ accepts.
This protocol uses $\Theta(n^{3})$ cards and $\Theta(kn^{2})$ shuffles.
@@ 5.2 Proof of Correctness and Security
  • In Step 4, since $P$ reconstructs a template that has just been used, all $n$ templates are correct, and thus this step will pass.
Therefore, every iteration will pass. After $k$ iterations, all integers in $P$’s extended solution will be printed on the grid, so all integers in the original puzzle will match the ones on the corresponding cards.
Hence, we can conclude that $V$ always accept. ∎
@@
Since Step 4 passes for every iteration, all $n$ templates are correct after each iteration (and also at the beginning of the protocol), which implies the integers printed in every iteration form a shape of a square and are equal to the side length of that square.
In Step 3, since the Tatami printer passes for every iteration, $B_{i}$ does not overlap with $B_{1},B_{2},...,B_{i-1}$ for every $i$. Also, since the final verification passes, the combined area of $B_{1},B_{2},...,B_{k}$ must cover the whole puzzle grid, i.e. $B_{1},B_{2},...,B_{k}$ form a partition of the grid. Furthermore, each point in the grid can be a corner point of at most two squares, so no four squares share a corner point.
@@ 6 Future Work
We developed the Tatami printer, which can check the corner constraint of Tatami puzzles. However, this protocol still has limitation that it cannot check other constraints involving a relation between multiple regions, e.g. a constraint in Fillmat which specifies that horizontally or vertically adjacent rectangles cannot have the same number. A possible future work is to find a technique to deal with these kinds of constraint.

### v2

@@ Abstract
Tatami puzzles are pencil puzzles with an objective to partition a rectangular grid into rectangular regions such that no four regions share a corner point, as well as satisfying other constraints. In this paper, we develop a physical card-based protocol called Tatami printer that can help verify solutions of Tatami puzzles. We then use the Tatami printer to construct zero-knowledge proof protocols for two such puzzles: Tatamibari and Square Jam. These protocols enable a prover to show a verifier the existence of the puzzles’ solutions without revealing them.
@@ 1 Introduction
Pencil puzzles are puzzles written on paper that can be solved using logical reasoning. Examples of them include Sudoku, Numberlink, and Nonogram. Pencil puzzles are classified into several categories based on their main theme. Puzzles involving partitioning a rectangular grid into multiple regions to satisfy certain rules are often called decomposition puzzles.
Many decomposition puzzles, especially the ones originated in Japan, specify that the regions must be rectangular, and also have a common corner constraint: no four regions can share a corner point. This rule came from the arrangement of Japanese Tatami mats. Hence, these puzzles can be collectively called Tatami puzzles, a subcategory of decomposition puzzles [11].
@@ 1.3 Zero-Knowledge Proof
Suppose Agnes designed a difficult Tatami puzzle and challenged her friend Brian to solve it. After a while, he could not solve the puzzle and began to doubt whether it actually has a solution. Agnes needs to convince him that her puzzle has a solution without revealing it (which would render the challenge pointless). A zero-knowledge proof (ZKP) makes this difficult task possible.
@@
In 1991, Goldreich et al. [2] proved that a ZKP exists for every NP problem, implying one can construct a computational ZKP for any NP pencil puzzle via a reduction. Such construction, however, is neither practical nor intuitive as it requires cryptographic primitives. Instead, many researchers focused on developing physical ZKPs using objects found in everyday life such as a deck of playing cards. These card-based protocols have the benefit that they do not require computers and also allow external observers to verify that the prover truthfully executes them (which is often a challenging task for digital protocols). They also have didactic values and can be used to teach the concept of ZKP to non-experts.
@@ 1.4 Card-based Zero-Knowledge Proof Protocols
There is a line of work dedicated to design physical card-based ZKP protocols for specific pencil puzzles, including ABC End View [14], Goishi Hiroi [14], Heyawake [12], Hitori [12], Juosan [9], Kakuro [10], Masyu [8], Nonogram [13], Numberlink [16], Nurikabe [12], Ripple Effect [17], Shikaku [15], Slitherlink [8], Sudoku [19], Sumplete [4], Takuzu [9], and Toichika [14].
In 2024, Ruangwises and Iwamoto [18] proposed a more generic protocol called printing protocol, which can help verify solutions of several decomposition puzzles such as Five Cells and Meadows. Their protocol, however, has a limitation that it cannot check constraints involving relationships between multiple regions, and thus cannot be used to verify solutions of Tatami puzzles.
@@ 1.5 Our Contribution
In this paper, we propose a card-based protocol called Tatami printer, which is modified from the printing protocol of Ruangwises and Iwamoto [18].
@@
We then use the Tatami printer to construct ZKP protocols for two Tatami puzzles: Tatamibari and Square Jam.
@@ 2.2 Pile-Shifting Shuffle
Given a matrix $M$ of cards, a pile-shifting shuffle [20] rearranges the columns of $M$ by a uniformly random cyclic shift. It can be implemented by putting all cards in each column into an envelope, and repeatedly picking some envelopes from the bottom and putting them on the top of the pile.
Each card in $M$ can be replaced by a stack of cards (as long as every stack in the same row has the same number of cards), and the protocol still works in the same way.
@@ 2.3 Chosen Cut Protocol
Each card in $C$ can be replaced by a stack of cards (as long as every stack has the same number of cards), and the protocol still works in the same way.
@@ 2.4 Printing Protocol of Ruangwises and Iwamoto
In a recent work, Ruangwises and Iwamoto [18] developed the following printing protocol.
A $p\times q$ matrix of face-down cards (called a template) and another $p\times q$ matrix of face-down cards of an area from the puzzle grid are given (the front side of all cards are known to $P$ but not to $V$). This protocol verifies that positions in the area corresponding to non-blank cards in the template are initially empty (consisting of all blank cards). Then, it places all non-blank cards from the template at the corresponding positions in the area, replacing the original blank cards (see Fig. 4) without revealing any card to $V$.
@@
Figure 4: Examples of a successful print from a $3\times 5$ template onto a $3\times 5$ area (top) and an unsuccessful print due to overlapping non-blank cards (bottom)
@@
      Reveal that the selected card is blank (otherwise $V$ rejects) and remove it from the stack.
After these steps, all non-blank cards from the template are placed at the corresponding positions in the area, and $V$ is convinced that these positions in the area were initially empty.
For decomposition puzzles with simple constraints, $P$ can repeatedly apply the printing protocol to print each region of the partition according to $P$’s solution onto the puzzle grid, convincing $V$ that these regions do not overlap.
However, a major limitation of this protocol is that each region is separately printed, so it cannot check constraints involving relationships between multiple regions. Therefore, it cannot be used to verify solutions of Tatami puzzles.
@@ 3 Our Proposed Protocol: Tatami Printer
Considering the limitation of the printing protocol, we aim to modify the protocol so that it can also check the corner constraint of Tatami puzzles.
@@
On top of each card in a template, we place a $\clubsuit$  if its top-left corner is a corner point of a rectangle the template represents, and a $\heartsuit$  if its top-left corner is not a corner point. Also, on top of each card in the puzzle grid, we place two cards (initially both $\heartsuit$ s) to keep track of the number of rectangles having its top-left corner as their corner point (which is the number of $\clubsuit$ s among them).
@@
When printing a $p\times q$ template onto a $p\times q$ area from the grid (the front side of all cards are known to $P$ but not to $V$), $P$ performs the following steps.
@@
      Reveal that the selected card is blank (otherwise $V$ rejects) and remove it from the stack.
@@
    For the corner counter part, place each card from the template on top of a corresponding two-card stack from the area, creating $pq$ three-card stacks.
@@
      Reveal that the selected card is a $\heartsuit$  (otherwise $V$ rejects) and remove it from the stack.
After these steps, all non-blank cards from the main part of the template are placed at the corresponding positions in the area. All $\clubsuit$ s from the corner counter part of the template are also added to the corresponding positions in the area. $V$ is convinced that these positions in the area were initially empty, and that each point in the area is a corner point of at most two rectangles.
@@ 4.1 Main Protocol
Suppose that in $P$’s extended solution, the grid is partitioned into $k$ rectangles $B_{1},B_{2},$ $,B_{k}$. Note that $k$ is public information, which is the number of cells containing a symbol in the original puzzle. For each $i=1,2,...,k$, $P$ performs the following steps.
@@
Finally, $P$ reveals all cards from the main part on the cells that contain a symbol in the original puzzle. $V$ verifies that the symbols on the cards match the ones in the cells (otherwise $V$ rejects). $P$ also reveals the main part of all dummy stacks that they are still blank (otherwise $V$ rejects). If all verification steps pass, then $V$ accepts.
@@ 4.2 Proof of Correctness and Security
Hence, we can conclude that $V$ always accepts. ∎
@@ 5 ZKP Protocol for Square Jam
However, unlike in Tatamibari, the number of squares in $P$’s solution is not public information, so $P$ cannot reveal that number to $V$. Therefore, $P$ has to apply the Tatami printer for $n^{2}$ times, sometimes printing nothing, as $V$ knows that there are at most $n^{2}$ squares.
In our protocol, $P$ constructs $n$ templates, one for each size of square. Each template has size $(n+1)\times(n+1)$, and the square is placed at the top-left corner of the template. A cell inside the square is represented by a card with a number equal to the side length of that square, while a cell outside the square is represented by a blank card (see Fig. 7). In addition, $P$ constructs a blank template of the same size, with the main part consisting of all blank cards and the corner counter part consisting of all $\heartsuit$ s.
@@ 5.1 Main Protocol
$P$ also constructs templates of all $n$ sizes of square, plus a blank template, and lets $V$ verify that all $n+1$ templates are correct (otherwise $V$ rejects).
Suppose that in $P$’s extended solution, the grid is partitioned into $k$ squares $B_{1},B_{2},...,B_{k}$. For each $i=1,2,...,n^{2}$, $P$ performs the following steps.
@@
    Apply the chosen cut protocol to select an $(n+1)\times(n+1)$ area whose top-left corner is the top-left corner of $B_{i}$ (if $i>k$, select any area).
@@
    Apply the chosen cut protocol to select a template of a square with the same size as $B_{i}$ (if $i>k$, select a blank template).
@@
    Reconstruct a template that has just been used and replenish the pile of templates with it. Let $V$ verify again that all $n+1$ templates are correct (otherwise $V$ rejects).
Finally, $P$ reveals all cards from the main part on the cells that contain an integer in the original puzzle. $V$ verifies that the integers on the cards match the ones in the cells (otherwise $V$ rejects). $P$ also reveals the main part of all dummy stacks that they are still blank (otherwise $V$ rejects). If all verification steps pass, then $V$ accepts.
This protocol uses $\Theta(n^{3})$ cards and $\Theta(n^{4})$ shuffles.
@@ 5.2 Proof of Correctness and Security
Suppose $P$ knows an extended solution of the puzzle. Consider each $i$-th iteration of the main protocol ($i\leq k$).
@@
  • In Step 4, since $P$ reconstructs a template that has just been used, all $n+1$ templates are correct, and thus this step will pass.
Therefore, the first $k$ iterations will pass. All subsequent iterations just print a blank template (we call them trivial iterations), so they will also pass. After the first $k$ iterations, all integers in $P$’s extended solution will be printed on the grid. Subsequent trivial iterations do not modify the grid, so all integers in the original puzzle will match the ones on the corresponding cards.
Hence, we can conclude that $V$ always accepts. ∎
@@
Since Step 4 passes for every iteration, all $n+1$ templates are correct after each iteration (and also at the beginning of the protocol), which implies the integers printed in every nontrivial iteration form a shape of a square and are equal to the side length of that square.
Let $D_{1},D_{2},...,D_{\ell}$ be the squares printed in each nontrivial iteration in this order, for some $\ell\leq n^{2}$. In Step 3, since the Tatami printer passes for every iteration, $D_{i}$ does not overlap with $D_{1},D_{2},...,D_{i-1}$ for every $i$. Also, since the final verification passes, the combined area of $D_{1},D_{2},...,D_{\ell}$ must cover the whole puzzle grid, i.e. $D_{1},D_{2},...,D_{\ell}$ form a partition of the grid. Furthermore, each point in the grid can be a corner point of at most two squares, so no four squares share a corner point.
@@ 6 Future Work
We developed the Tatami printer protocol, which can check the corner constraint of Tatami puzzles. However, this protocol still has a limitation that it cannot check other types of constraint involving relationships between multiple regions, e.g. a constraint in Fillmat which specifies that horizontally or vertically adjacent rectangles cannot contain the same number. An interesting future work is to develop a technique to check such constraints.